OpenStudy (anonymous):

A 0.3-kg ball has a velocity of 30 m/s. (a) What is the kinetic energy of the ball? (b) How much work would be required to stop the ball?

OpenStudy (anonymous):

\[KE = 1/2mv^2\] \[W = F*dcos(\theta)\]

OpenStudy (anonymous):

where d is the displacement